# In the Network of the Conclave: Social Network Analysis and the Making of a Pope

### v1

@@ 4.1 Data Collection
To investigate the network mechanisms that permeate the conclave, we collected a broad range of data on current members of the College of Cardinals from publicly available sources. Data collection was conducted following two different logics. The first approach was informed by the assumption that the social structure connecting cardinals possesses a dual nature: it reflects both the current organizational configuration and historically constructed relationships whose influence persists over time. In addition to that, participation in collegial bodies cannot be equated with membership in a peer group, because the various Roman Curia entities that we investigated include hierarchical roles, such as secretary and under-secretary. Thus, we conducted an initial study focusing exclusively on the relationships among cardinals occupying apical positions, while adopting a historical perspective that could also capture past co-membership ties among them within these collegial bodies. To construct the overall network, this affiliation network was subsequently overlaid with another network derived from consecration relationships. Additionally, we also included informal relationships among cardinals, which were identified through reputable journalistic articles. [^7] E.g., https://www.catholicculture.org; https://www.pillarcatholic.com; https://www.indcatholicnews.com; https://angelusnews.com/. As mentioned, findings from this first study were disseminated to a broad audience via social media and newspaper publications before the election of Cardinal Prevost as Pope Leo XIV. For completeness, results and graphs from this data collection are fully reported in the Appendix of this article. It is important to note that the results of this first analyses and those of the more comprehensive investigation presented in this article are generally convergent in identifying key network figures.
The second approach and study, which is the focus of this paper, involved the reconstruction of a broader social structure encompassing all members of the collegial bodies, not just apical roles, as of the date of the data collection period (April 2025). Specifically, we collected data through official Vatican archives, which we used to obtain information regarding the composition of the collegial bodies of the Roman Curia (e.g., commissions, dicasteries, etc…), and thus the institutional affiliations of all cardinals involved with these units. [^8] https://press.vatican.va/
@@
Furthermore, to ensure the accuracy of the data regarding the membership composition, we checked the affiliation of each cardinal by consulting their own biographical profiles, [^9] https://press.vatican.va/content/salastampa/en/documentation/card_bio_typed.html which record their active memberships in the curial bodies. We did so for both voting and non-voting cardinals. Despite these careful checks, however, it is possible that a small number of ties may be missing due to official sources not being fully up to date or listing affiliations across multiple, inconsistently structured sections. However, we have reasons to believe that such instances are likely to represent a minority of cases. Moreover, prior research in social network analysis has shown that centrality measures tend to be robust to small amounts of missing data (Borgatti et al.,, 2006). As such, we do not expect these potential omissions to meaningfully affect our results. For the purposes of this study, we limited our analysis to active organizational affiliations in the collegial bodies as of the extraction date (April 2025). By anchoring our dataset to a specific date, we aimed to construct a balanced and consistent representation of the curial landscape at the time of the conclave. Such a standardizing of the temporal frame and sourcing methods also facilitates meaningful comparisons with future research on the College of Cardinals.
@@
Finally, we also collected additional information on each cardinal’s age (as of April 2025) and nationality using the same website. To support our visualizations and better contextualize the network, we incorporated data on each voting cardinal’s doctrinal orientation regarding key pastoral issues. These include positions on topics such as the blessing of same-sex couples, the ordination of female deacons, communion for divorced and remarried Catholics, and the decentralization of Church authority. These stances, which can be drawn from public statements, writings, and appointments, help illustrate underlying doctrinal alignments within the College of Cardinals. In our case, this information was obtained from an analysis published by The Sunday Times, which classified cardinals based on their public statements along a spectrum ranging from strong liberal to strong conservative. [^11] https://www.thetimes.com/uk/religion/article/inside-papal-conclave-politics-wjb9dhdvs [^12] All data we described is accessible at the following link: https://github.com/leonardorizzo/Conclave2025
@@ 4.2 Data Processing
We begin our analysis by constructing a two-mode (i.e., bipartite) network that captures all affiliations between cardinals and the various collegial bodies of the Roman Curia, allowing for a clear visualization of institutional co-memberships. Figure 1 visualizes this network, which includes 160 nodes and 352 edges.
@@
We then integrated the formal co-membership network with the co-consecration network, resulting in a multiplex network comprising 188 cardinals and 2,884 weighted edges. Figures 4 and 5 shows the resulting networks. In particular, while Figure 4 leverages a standard spring-embedding layout algorithm, Figure 5 displays a block model representation with edge bundling, which enhances the readability of the network by visually emphasizing cluster structures and intergroup connections (Holten,, 2006).
@@ 5 Network Analysis
To identify the structural foundations of potential influence within the conclave, we analyzed the network positions occupied by cardinals in the College of Cardinals. As mentioned, the network we analyzed embodies both co-membership ties, which arise from joint service in collegial bodies, and consecration ties, reflecting apostolic lineage. We examined the three mechanisms introduced earlier—status, mediation power, and coalition building—and operationalized each of them using well-established network indicators. For status, we used eigenvector centrality to capture the extent to which a cardinal is well connected to other central cardinals (Podolny,, 2005). For mediation power, we computed betweenness centrality as a proxy for a cardinal’s capacity to bridge across disconnected parts of the network (Freeman,, 1977). For coalition building, we computed a composite index that integrates three key parameters: (1) community density, measured within each cardinal’s modularity-based community as identified via modularity maximization (Reichardt and Bornholdt,, 2006); (2) the size of the egonetwork, reflecting the number of direct connections held by the cardinal; and (3) the inter-community reach, calculated as the proportion of distinct communities connected through direct ties relative to the total number of communities in the network.
@@ 5.1 Status
Eigenvector centrality captures not only how many connections an actor has, but also how well-connected their neighbors are, thus making it especially suited for identifying those who are embedded within the core of the network’s power structure (Bonacich,, 2007). In this context, a high eigenvector score suggests that a cardinal is not only prominent, but connected to other prominent actors, and thus enjoys high structural visibility. The ranking in Table 1 below shows the top 15 cardinals by weighted eigenvector centrality. These individuals occupy structurally privileged positions that may afford them greater informal influence, visibility, and endorsement potential within the conclave.
@@
As the table shows, Cardinals Prevost, Tagle, and Gugerotti emerge as the most centrally embedded actors in terms of eigenvector centrality, followed closely by Fernández and Parolin, who at the time of the conclave held the key formal role of Secretary of State. Their high centrality scores reflect not only extensive ties within the College, but also the strategic quality of those ties, positioning them at the core of ecclesiastical network. These cardinals may be especially well-placed to attract early attention, mobilize endorsements, or signal institutional continuity within the deliberative dynamics of the conclave.
@@ 5.2 Mediation Power
We calculated each cardinal’s betweenness centrality on the combined network of co-membership and consecration ties. Betweenness centrality measures how frequently a node lies on the shortest paths between all other pairs of nodes (Freeman,, 1977), capturing an actor’s capacity to connect disparate parts of the network. We use the weighted version of this metric, which accounts for the strength of ties (rather than treating all edges as equal). Weighted betweenness is particularly useful in multi-layered networks, where stronger ties may play a more significant role in shaping real-world interactions. Table 2 presents the top cardinals ranked by weighted betweenness centrality:
@@
Cardinals Betori, Tobin, and Zuppi emerge as powerful brokers. Figures like Cardinal Czerny and Cardinal De Donatis also appear prominently, reflecting their role in connecting across different communities within the College of Cardinals. Interestingly, some cardinals such as Prevost, Tagle, and Koch who ranked highly on eigenvector centrality (status), also appear here—suggesting a dual advantage: they are both visible within central clusters and capable of brokering across subgroups. This convergence of structural prominence and brokerage potential may afford them particular influence in informal consensus-building processes, even if not all high-betweenness cardinals are considered papabili in public discourse.
@@ 5.3 Coalition Building
The Coalition Building Index is calculated as a weighted average of the three components, each assigned a specific weight. Community cohesion is defined as the internal density of the cardinal’s community, detected using a modularity-based algorithm. Egonetwork size is measured by the cardinal’s number of direct connections (degree), normalized by the maximum degree in the network. Cross-community reach captures the share of other communities to which a cardinal is directly connected. Together, these elements provide a replicable and theory-informed measure of a cardinal’s coalition-building capacity based solely on their network position. Table 3 below presents the top cardinals by coalition-building index:
@@ 5.4 Underdogs and Kingmakers: The Network Advantage of Cardinal Prevost
When it comes to forecasting frontrunners, many tend to favor visible markers of influence: tenure, title, charisma. The papabili lists that circulated ahead of the 2025 conclave were no exception. Pundits, commentators, and bookmakers clustered around highly visible members of the Roman Curia and prominent archbishops, that is, those with formal authority, diplomatic profiles, or public recognition.
@@
The contrast could not be sharper. While public forecasts focused on surface-level attributes, our structural indicators captured invisible, yet consequential, markers of influence. Prevost was not the loudest voice in the room, but he was the one most likely to connect the right people at the right moment. The textitpapabili lists built on visible credentials missed what the network revealed. And in the end, the network got it right. The election of Cardinal Prevost—Pope Leo XIV—was not just the result of personal qualities. It was also the result of a strong structural position. We represent this structural position in Figure 6, which features Cardinal Prevost’s egonetwork. The cardinals positioned in a circle around him represent his direct connections. At the center, the pie chart illustrates Prevost’s affiliations across ten collegial bodies. Each surrounding node also features a pie chart, indicating that cardinal’s co-memberships with Prevost in those same bodies. The greater the number of shared affiliations (i.e., the more colored segments in the pie), the stronger and more multiplex the connection between Prevost and that cardinal.
INSERT FIGURE 6 ABOUT HERE
Our model also highlights another similar case to Cardinal Prevost: Cardinal José Tolentino de Mendonça. Despite occupying a structurally prominent position within the network in some structural dimensions, he was largely absent from public papabili lists. One likely reason is his age: at 59, he may have been perceived as relatively young to be considered a serious contender. Yet it is worth noting that he is not significantly younger than Cardinal Pizzaballa, who appeared in many predictive models. This suggests that Tolentino’s low visibility was less a matter of eligibility and more a reflection of how traditional analyses continue to overlook structural indicators of influence.
5.5 Robustness Analysis
@@
Additionally, we tested the impact of incorporating the informal ties previously identified through journalistic sources. Due to their limited number, these ties had no significant effect on the overall network structure or rankings. Similarly, substituting unweighted measures for eigenvector centrality did not alter the cardinal ranking for the Status metric. However, an interesting result emerged when applying the unweighted version of betweenness centrality: Cardinal Prevost rose to the second-highest position within the Conclave in terms of mediation power, highlighting his strategic location in the network.
Finally, we conducted a focused analysis using only the formal institutional network, excluding all consecration ties. Under this configuration, Cardinal Prevost’s network centrality increased significantly across multiple dimensions. He ranked first in status, third in mediation power and first in coalition-building capacity, underscoring his dominant structural position within the formal curial network. This result highlights Prevost’s strategic advantage in the realm of institutional affiliations, even in the absence of ecclesiastical lineage ties.
@@ 6 Discussion
Our reconstruction of the College of Cardinals as a multiplex network offers a powerful lens through which to explore the sacred mysteries of the conclave: existing alongside the formal hierarchical structure, this network provides cardinals with both opportunities and constraints in mobilizing critical resources consequential to the decision-making process. The conclave may be wrapped in ritual, but influence still follows relational pathways that are as patterned, and as measurable, as those in any corporate boardroom or parliamentary council (Westphal and Zajac,, 1997). By foregrounding structure, we reaffirm a foundational insight of network research: position is a structural signature of influence.
As we discussed, the practical payoff of this lens is summarized by how Cardinal Prevost became Pope Leo XIV. On the eve of the conclave most bookmakers [^14] Following a CNBC report, bettors poured more than \$40 million into the papal conclave, on just two prediction market platforms: over \$30 million on Polymarket, according to its website, and over \$10.6 million on Kalshi, says a company spokesperson. Cardinal Prevost had odds of less than 1%, according to Kalshi’s website. Of the over 33,000 conclave trades placed on Kalshi, just 416 trades totaling to nearly \$450,000 were placed on Prevost. and pundits relegated him to the margins, dazzled instead by more visible media favorites. A network reading told a different story. The structural visibility and informational leverage that flowed from that dual position placed Prevost squarely within the set of credible frontrunners. That conventional commentary overlooked these facts is not a failure of information; it is a failure of perspective. When analysis treats individuals as isolated units, structural advantage becomes invisible.
@@
Our study does not provide sufficient elements to offer an answer to the question of which individual and organizational factors led Cardinal Prevost to occupy such a prominent position in the network. Yet, we speculate that this was anything but a random process. Cardinal Prevost’s trajectory within the Roman Curia was notably accelerated under Pope Francis, who entrusted him with a series of increasingly influential roles. Beginning with his appointment as Apostolic Administrator of Chiclayo in 2014, Prevost was subsequently named bishop of the same diocese in 2015, and progressively took on responsibilities both in Peru and within the universal Church. His appointments to multiple key dicasteries—including the key Dicastery for Bishops and the Congregation for the Clergy—culminated in his 2023 nomination as Prefect of the Dicastery for Bishops (which has the important task of selecting and supervising bishops) and President of the Pontifical Commission for Latin America, all positions of substantial influence in the governance of the Church. Created cardinal in September 2023, he entered the order of cardinal bishops in February 2025, receiving the suburbicarian see of Albano. To what extent all these appointments—which enhanced Cardinal Prevost’s prominence e visibility—were the result of an intentional design by Pope Francis and part of a plausible “succession plan” is impossible to determine. However, it is certainly a question that warrants further study. If the answer were affirmative, it would suggest that in a highly complex organization, the composition of collegial bodies—much more than formal hierarchy—facilitates the formation of networks that are highly relevant to key decision-making processes and that networks within organizations tend to be draped around the scaffolding of formal arrangements (Soda,, 2025). Initiatives aimed at shaping interpersonal relationships and the structures of intra-organizational networks could be effectively anchored in modifications to the formal processes of the organization (Ouchi and Dowling,, 1974). Although these interventions are initiated through formal changes—such as alterations to structures, processes, and organizational systems—their effects on social structures can be systematically analyzed using network analysis methodologies (Valente,, 2012).
@@
Our study on the conclave also offers a cautionary note for the current infatuation with big data. Big data without theory is the statistical equivalent of a telescope without a lens: longer range, perhaps, but no clarity. Predictive dashboards and AI pipelines that scrape speeches, tweets, or betting markets missed Cardinal Prevost because they modelled attributes and signals, while forgetting relationships (Antonioni et al.,, 2025). The implication is not that machine learning is futile, but that strategic value lies in asking the right questions and using the right theoretical lenses. In this study, we argued that social network analysis is a powerful tool to investigate the conclave.

### v2

@@ 4.1 Data Collection
To investigate the network mechanisms that permeate the conclave, we collected a broad range of data on current members of the College of Cardinals from publicly available sources. Data collection was conducted following two different logics. The first approach was informed by the assumption that the social structure connecting cardinals possesses a dual nature: it reflects both the current organizational configuration and historically constructed relationships whose influence persists over time. In addition to that, participation in collegial bodies cannot be equated with membership in a peer group, because the various Roman Curia entities that we investigated include hierarchical roles, such as secretary and under-secretary. Thus, we conducted an initial study focusing exclusively on the relationships among cardinals occupying top positions, while adopting a historical perspective that could also capture past co-membership ties among them within these collegial bodies. To construct the overall network, this affiliation network was subsequently overlaid with another network derived from consecration relationships. Additionally, we also included informal relationships among cardinals, which were identified through reputable journalistic articles. [^7] https://www.catholicculture.org; https://www.pillarcatholic.com; https://www.indcatholicnews.com; https://angelusnews.com/. As mentioned, findings from this first study were disseminated to a broad audience via social media and newspaper publications before the election of Cardinal Prevost as Pope Leo XIV. For completeness, results and graphs from this data collection are fully reported in the Appendix of this article. It is important to note that the results of this first analyses and those of the more comprehensive investigation presented in this article are generally convergent in identifying key network figures.
@@
The second approach and study, which is the focus of this paper, involved the reconstruction of a broader social structure encompassing all members of the collegial bodies, not just top roles, as of the date of the data collection period (April 2025). Specifically, we collected data through official Vatican archives, which we used to obtain information regarding the composition of the collegial bodies of the Roman Curia (e.g., commissions, dicasteries, etc…), and thus the institutional affiliations of all cardinals involved with these units. [^8] https://press.vatican.va/ Furthermore, to ensure the accuracy of the data regarding the membership composition, we checked the affiliation of each cardinal by consulting their own biographical profiles, [^9] https://press.vatican.va/content/salastampa/en/documentation/card_bio_typed.html which record their active memberships in the curial bodies. We did so for both voting and non-voting cardinals. Despite these careful checks, however, it is possible that a small number of ties may be missing due to official sources not being fully up to date or listing affiliations across multiple, inconsistently structured sections. However, we have reasons to believe that such instances are likely to represent a minority of cases. Moreover, prior research in social network analysis has shown that centrality measures tend to be robust to small amounts of missing data (Borgatti et al.,, 2006). As such, we do not expect these potential omissions to meaningfully affect our results. For the purposes of this study, we limited our analysis to active organizational affiliations in the collegial bodies as of the extraction date (29 April 2025). By anchoring our dataset to a specific date, we aimed to construct a balanced and consistent representation of the curial landscape at the time of the conclave. Such a standardizing of the temporal frame and sourcing methods also facilitates meaningful comparisons with future research on the College of Cardinals.
@@
Finally, we also collected additional information on each cardinal’s age (as of April 2025) and nationality using the same website. To support our visualizations and better contextualize the network, we incorporated data on each voting cardinal’s doctrinal orientation regarding key pastoral issues. These include positions on topics such as the blessing of same-sex couples, the ordination of female deacons, communion for divorced and remarried Catholics, and the decentralization of Church authority. These stances, which can be drawn from public statements, writings, and appointments, help illustrate underlying doctrinal alignments within the College of Cardinals. In our case, this information was obtained from an analysis published by The Sunday Times, which classified cardinals based on their public statements along a spectrum ranging from strong liberal to strong conservative. [^11] https://www.thetimes.com/uk/religion/article/inside-papal-conclave-politics-wjb9dhdvs, [^12] All data we described is accessible at the following link: https://github.com/leonardorizzo/Conclave2025
@@ 4.2 Data Processing
We begin our analysis by constructing a two-mode (i.e., bipartite) network that captures all affiliations between cardinals and the various collegial bodies of the Roman Curia, allowing for a clear visualization of institutional co-memberships. Figure 1 visualizes this network, which includes 160 nodes (136 cardinals and 24 entities) and 352 edges.
@@
We then integrated the formal co-membership network with the co-consecration network, resulting in a multiplex network comprising 188 cardinals and 2,884 weighted edges. Table 1 summarizes the key network statistics and associated data sources. Figures 4 and 5 shows the resulting networks. In particular, while Figure 4 leverages a standard spring-embedding layout algorithm, Figure 5 presents a community-based visualization of the network using edge bundling, which improves readability by highlighting intra-community cohesion and inter-community linkages (Holten,, 2006).
INSERT TABLE 1 ABOUT HERE
@@ 5 Network Analysis
To identify the structural foundations of potential influence within the conclave, we analyzed the network positions occupied by cardinals in the College of Cardinals. As mentioned, the network we analyzed embodies both co-membership ties, which arise from joint service in collegial bodies, and consecration ties, reflecting apostolic lineage. We examined the three mechanisms introduced earlier—status, mediation power, and coalition building—and operationalized each of them using well-established network indicators. [^13] Centrality measures were computed on the unweighted network. Robustness checks using weighted ties yielded consistent results (Section 5.5). For status, we used eigenvector centrality to capture the extent to which a cardinal is well connected to other central cardinals (Podolny,, 2005). For mediation power, we computed betweenness centrality as a proxy for a cardinal’s capacity to bridge across disconnected parts of the network (Freeman,, 1977). For coalition building, we computed a composite index that integrates three key parameters: community density, the size of the egonetwork, and the inter-community reach.
@@ 5.1 Status
Eigenvector centrality captures not only how many connections an actor has, but also how well-connected their neighbors are, thus making it especially suited for identifying those who are embedded within the core of the network’s power structure (Bonacich,, 2007; Aslarus et al.,, 2025). In this context, a high eigenvector score suggests that a cardinal is not only prominent, but connected to other prominent actors, and thus enjoys high structural visibility. The ranking in Table 2 below shows the top 15 cardinals by eigenvector centrality. These individuals occupy structurally privileged positions that may afford them greater informal influence, visibility, and endorsement potential within the conclave.
@@
As the table shows, Cardinals Prevost, Tagle, and Parolin, who at the time of the conclave held the key formal role of Secretary of State, emerge as the most centrally embedded actors in terms of eigenvector centrality, followed closely by Cardinals Gugerotti and Fernández. Their high centrality scores reflect not only extensive ties within the College, but also the strategic quality of those ties, positioning them at the core of ecclesiastical network. These cardinals may be especially well-placed to attract early attention, mobilize endorsements, or signal institutional continuity within the deliberative dynamics of the conclave.
@@ 5.2 Mediation Power
We calculated each cardinal’s betweenness centrality on the combined network of co-membership and consecration ties. Betweenness centrality measures how frequently a node lies on the shortest paths between all other pairs of nodes (Freeman,, 1977), capturing an actor’s capacity to connect disparate parts of the network. Table 3 presents the top cardinals ranked by betweenness centrality:
@@
Cardinals Tagle, Prevost, and Betori emerge as powerful brokers. Figures like Cardinal Fernández and Cardinal Tobin also appear prominently, reflecting their role in connecting across different communities within the College of Cardinals. Interestingly, some cardinals such as Prevost, Tagle, Parolin, and Fernández who ranked highly on eigenvector centrality (status), also appear here—suggesting a dual advantage: they are both visible within central clusters and capable of brokering across subgroups. This convergence of structural prominence and brokerage potential may afford them particular influence in informal consensus-building processes, even if not all high-betweenness cardinals are considered papabili in public discourse.
@@ 5.3 Coalition Building
The first step involves the identification of communities within the network. Community detection seeks to locate densely connected regions—subsets of nodes that are more strongly interconnected relative to the rest of the network—and assign them to the same group. While numerous algorithms exist for this purpose, many are stochastic in nature and require the calibration of one or more parameters. To ensure reproducibility and reduce methodological complexity, we employed a deterministic approach based on greedy modularity optimization (Clauset et al.,, 2004). [^15] The algorithm identified 10 communities within the connected component, with a modularity score of 0.19. It is important to note that the metrics presented in the subsequent analysis are highly sensitive to the underlying community structure, and may vary substantially depending on the chosen detection method.
The Coalition Building Index is calculated as a weighted average of the three components, each assigned a specific weight. Community cohesion is defined as the internal density of the cardinal’s community, expressed as a value between 0 and 1. Egonetwork size is measured by the cardinal’s number of direct connections (degree), normalized by the maximum degree in the network. Cross-community reach captures the share of other communities to which a cardinal is directly connected; this value also ranges between 0 and 1. Together, these elements provide a replicable and theory-informed measure of a cardinal’s coalition-building capacity based solely on their network position. Table 4 below presents the top cardinals by coalition-building index:
@@ 5.4 Underdogs and Kingmakers: The Network Advantage of Cardinal Prevost
When it comes to forecasting frontrunners, many tend to favor visible markers of influence: tenure, title, charisma. The papabili lists that circulated ahead of the 2025 conclave were no exception. With very few exceptions, pundits, commentators, and bookmakers clustered around highly visible members of the Roman Curia and prominent archbishops, that is, those with formal authority, diplomatic profiles, or public recognition
@@
This contrast is striking. While public forecasts emphasized surface-level attributes, our structural indicators captured less visible but meaningful markers of influence. Prevost was not the most prominent figure publicly, yet he demonstrated a high likelihood of connecting the right individuals at critical moments. The papabili lists, grounded in visible credentials, overlooked insights revealed through the network. The election of Cardinal Prevost—Pope Leo XIV—was not just the result of personal qualities. It was also the result of a strong structural position.
Our model also highlights another similar case to Cardinal Prevost: Cardinal José Tolentino de Mendonça. Despite occupying a structurally prominent position within the network in some structural dimensions, he was largely absent from public papabili lists. One likely reason is his age: at 59, he may have been perceived as relatively young to be considered a serious contender. Yet it is worth noting that he is not significantly younger than Cardinal Pizzaballa, who appeared in many predictive models. This suggests that Cardinal Tolentino’s low visibility was less a matter of eligibility and more a reflection of how traditional analyses continue to overlook structural indicators of influence.
5.5 Robustness Analyses
@@
Additionally, we tested the impact of incorporating the informal ties previously identified through journalistic sources. Due to their limited number, these ties had no significant effect on the overall network structure or rankings. Likewise, substituting weighted variants of centrality measures—accounting for the strength of ties—for eigenvector and betweenness centrality did not yield substantial changes in the cardinal rankings.
We also examined both weighted and unweighted degree centrality as alternative metrics and found consistent results: Cardinal Prevost ranked first, further reinforcing his prominence across different dimensions of network influence. In addition, we computed a measure of 2-step local eigenvector centrality to address a potential limitation of standard eigenvector centrality, which can assign disproportionately high scores to nodes embedded in large cliques. By focusing on influence within a two-step neighborhood, this measure reduces the inflationary effect of cliquishness. Even with this approach Cardinal Prevost remained at the top of the status ranking.
Finally, we conducted a focused analysis using only the formal institutional network, excluding all consecration ties. Under this configuration, Cardinal Prevost’s network centrality increased significantly across multiple dimensions. He ranked first across all of the three dimensions, underscoring his dominant structural position within the formal curial network. This result highlights Prevost’s strategic advantage in the realm of institutional affiliations, even in the absence of ecclesiastical lineage ties.
@@ 6 Discussion
Our reconstruction of the College of Cardinals as a multiplex network offers a powerful lens through which to explore the complex dynamics of the conclave: existing alongside the formal hierarchical structure, this network provides cardinals with both opportunities and constraints in mobilizing critical resources consequential to the decision-making process. The conclave may be wrapped in ritual, but influence still follows relational pathways that are as patterned, and as measurable, as those in any corporate boardroom or parliamentary council (Westphal and Zajac,, 1997). By foregrounding structure, we reaffirm a foundational insight of network research: position is a structural signature of influence.
As we discussed, the practical payoff of this lens is summarized by how Cardinal Prevost became Pope Leo XIV. On the eve of the conclave most bookmakers [^16] Following a CNBC report, bettors poured more than \$40 million into the papal conclave, on just two prediction market platforms: over \$30 million on Polymarket, according to its website, and over \$10.6 million on Kalshi, says a company spokesperson. Cardinal Prevost had odds of less than 1%, according to Kalshi’s website. Of the over 33,000 conclave trades placed on Kalshi, just 416 trades totaling to nearly \$450,000 were placed on Prevost. and pundits relegated him to the margins, overshadowed instead by more visible media favorites. The visibility of these others candidates was driven by individual attributes that were easily observable from the outside. For instance, Cardinal Parolin was considered a frontrunner due to his formal organizational rank as Secretary of State. Cardinal Tagle, by contrast, was recognized by public opinion for his extroverted style, participatory leadership, and attentiveness to youth issues. Cardinal Besungu, in addition to his extensive experience, was seen as the most plausible candidate from the African continent, which has never in its history produced a pope. Cardinal Dolan was viewed as papabile in the event that a conservative candidacy prevailed in the conclave, while Cardinal Zuppi was among the leading figures associated with the progressive camp. The network perspective we adopted offers a different angle to understand the outcome of such a complex decision-making process. Specifically, the roles played within the co-membership network, which are visible within the College of Cardinals, and the informational leverage derived from his structural position placed Prevost squarely among the set of credible frontrunners.
The tension between attributes and relationships in shaping social outcomes is particularly salient in the context of the papal conclave, where it becomes visible in the contrast between prevailing media narratives and our network analysis. Public commentary has largely focused on observable individual characteristics—such as age, nationality, theological orientation or the formal role within the Roman Church—as key indicators of papal viability. In contrast, our approach highlights structural embeddedness as a more powerful, albeit less visible to the outside, marker of influence. This divergence recalls the foundational argument of Wellman and Berkowitz, (1988), who emphasized how behavior is often better explained in terms of relations rather than attributes, while also acknowledging their interaction.
Our analysis echoes the idea that network position is a robust correlate of behaviors, including voting decision as in the conclave. Yet we do not claim that structural advantage exists independently of individual traits or agency. A cardinal known for his doctrinal and political moderation, for instance, may be more likely to span ideological divides, suggesting that centrism operates not only as a personal attribute but also as a facilitator of bridging ties. This entanglement complicates causal inference: does influence stem from theological stance, from structural position, or from the mutually reinforcing alignment between the two? In other words, the relationship between personal characteristics and structural embeddedness raises important questions of endogeneity. From a theoretical perspective, several mechanisms may plausibly connect cardinals’ attributes, their structural positions, and their likelihood of receiving votes. In an additive logic, attributes and network positions may have an independent and direct effect on outcomes (Mehra et al.,, 2001). Alternatively, attributes may shape network structure—self-monitoring, for example, may facilitate cross-cutting ties (Casciaro et al.,, 2015; Sasovova et al.,, 2010)—thus indirectly influencing vote outcomes through relationships. Conversely, a cardinal’s central position in the network may shape his leadership skills or increase the perceived legitimacy of his traits, making structure an antecedent of malleable traits and skills. Furthermore, individual attributes and social structures could interact to explain behaviors. This idea is in line with “interaction models” of network which suggest how outcomes are shaped by a complementarity effect between structural network positions and individual traits (Soda et al.,, 2018).
In this discussion, however, it is important to consider the nature of ties which represents an important boundary condition. We study a context where ties among the cardinals emanate from a co-membership networks based on institutionalized patterns of shared committee service and episcopal lineages. These ties are formally mandated and thus relatively stable. Unlike informal interpersonal ties, these connections are not easily initiated or dissolved through individual agency, at least in the short term. As such, our networks are less plastic and less prone to ego-centric manipulation (Tasselli and Kilduff,, 2021).
@@
As the networks examined are the result of organizational design rather than the outcome of a self-organizing process, the role of agency and individual factors is attenuated or at least unfolding through actions aimed at influencing organizational decision-makers (Westphal and Stern,, 2007). It was the organizational decisions regarding the composition of collegial bodies that determined Prevost’s structural prominence. In this respect, Cardinal Prevost’s career within is particularly revealing. His trajectory within the Roman Curia was notably accelerated under Pope Francis, who entrusted him with a series of increasingly influential roles. Beginning with his appointment as Apostolic Administrator of Chiclayo in 2014, Prevost was subsequently named bishop of the same diocese in 2015, and progressively took on responsibilities both in Peru and within the universal Church. His appointments to multiple key dicasteries—including the key Dicastery for Bishops and the Congregation for the Clergy—culminated in his 2023 nomination as Prefect of the Dicastery for Bishops (which has the important task of selecting and supervising bishops) and President of the Pontifical Commission for Latin America, all positions of substantial influence in the governance of the Church. Created cardinal in September 2023, he entered the order of cardinal bishops in February 2025, receiving the suburbicarian see of Albano. To what extent all these appointments, which enhanced Cardinal Prevost’s prominence and visibility, were the result of an intentional design by Pope Francis and part of a plausible “succession plan” is impossible to determine.
In sum, what we have uncovered it that in a highly complex organization, the composition of collegial bodies—much more than formal hierarchy—facilitates the formation of networks that are highly relevant to key decision-making processes and that networks within organizations tend to be draped around the scaffolding of formal arrangements (Soda,, 2025).
Contributing to organizational network research, our study highlights the relevance of social infrastructures designed by organizational choices that foster connections among members beyond hierarchical ties. In the case of the conclave, the relational space generated through co-participation in collegial bodies played an important role in shaping voting behavior and consensus-building processes. More broadly, our study underscores the importance of accurately reconstructing all relational spaces—both formal and informal—that emerge within organizations, in order to better understand their underlying dynamics.
From a network intervention standpoint, initiatives aimed at shaping interpersonal relationships and the structures of intra-organizational networks could be effectively anchored in design choices (Ouchi and Dowling,, 1974). Although these interventions are initiated through formal changes—such as alterations to structures, processes, and organizational systems—their effects on social structures can be systematically analyzed using network analysis methodologies (Valente,, 2012).
@@
Another implication of our study is that it offers a cautionary note for the current infatuation with big data and the use of AI. Predictive dashboards and AI pipelines that scrape speeches, tweets, or betting markets missed Cardinal Prevost because they modelled attributes and signals, while forgetting relationships (Antonioni et al.,, 2025). The implication is not that machine learning is futile, but that strategic value lies in asking the right questions and using the right theoretical lenses. In this study, we argued that social network analysis is a powerful tool to investigate the conclave. Needless to say, this approach is deeply rooted in a long tradition of studies and research. From Katz,’s (1957) diffusion thresholds to Freeman,’s (1977) betweenness to Granovetter (1973) weak ties, the literature on networks has long showed that position matters. The conclave reminds us that classical tools remain potent when applied with care. By projecting a bipartite affiliation structure and layering a sacramental lineage onto it, we have shown how simple techniques illuminate a domain that, at first glance, seems empirically intractable.